OpenStudy (anonymous):

A square pyramid is 8 m on each side. The surface area is 240 m2. What is the slant height?

OpenStudy (anonymous):

Please Help :( Im really confused

OpenStudy (anonymous):

Well I'm a little confused too but Ima throw some things that I see and hopefully we can find something out.

OpenStudy (anonymous):

We know the area of the square is 8*8 right? 8*8=64

OpenStudy (anonymous):

Yes thats true...

OpenStudy (anonymous):

if we subtract that from the total surface area we get 240-64=176

OpenStudy (anonymous):

The pyramid will create 4 slant surfaces divide that evenly we get 176/4=44

OpenStudy (anonymous):

Hmmm I didnt think of it that way...

OpenStudy (anonymous):



OpenStudy (anonymous):

we said the area is 44

OpenStudy (anonymous):

we also know that we can find the area of the triangle by \[\frac{1}{2}*b*h\]

OpenStudy (anonymous):

where the base=8

OpenStudy (anonymous):

so 4*h=44

OpenStudy (anonymous):

h=11

OpenStudy (anonymous):